# Bridging Semantic Understanding and Popularity Bias with LLMs

### v1

@@ 1. Introduction
As recommender systems strive to become more trustworthy, enhancing semantic understanding of core echical concepts, such as popularity bias, is essential for advancing cognitive capabilities of web applications (Zhang et al., 2025b; Gao et al., 2025). Popularity bias occurs when popular items are recommended too frequently, while other products that may deserve more attention are ignored (Abdollahpouri et al., 2019). This surface-level issue, however, masks a more complex interaction between user preferences, item rankings, and recommendation algorithms (Tang et al., 2024). In this context, semantic understanding refers to a model’s ability to grasp the deeper implications of popularity bias, rather than merely recognizing its surface-level features (Xu, 2025). In other words, a comprehensive semantic understanding of popularity bias requires recognizing not only the observable imbalance in recommendations but also the underlying causal mechanisms that give rise to it.
@@
Developing a deeper semantic understanding of popularity bias is key to addressing challenges such as the long-tail problem (Zhang et al., 2025a), product homogeneity (Lin et al., 2025), and personalized user satisfaction (Silva et al., 2021; Zhang et al., 2025c). Existing methods, such as PAAC (Cai et al., 2024), primarily focus on surface-level mitigation by increasing recommendation diversity or boosting exposure to non-popular items. Although terms like “diversity” and “debiasing” are frequently used in prompts, our investigation with Large Language Models (LLMs) reveals that such cues remain shallow representations of popularity bias (Klimashevskaia et al., 2024; Xiang et al., 2025). They fail to encode the causal semantics that link user behavior, item features, and system design to the emergence of bias.
@@ 2.1. Semantic Understanding for Recommender Systems
As web applications continue to evolve, achieving a deeper semantic understanding of complex definitions is crucial for enhancing the cognitive capabilities of models (Zhang et al., 2024a, 2025d). Bhattacharya and Pandey (Bhattacharya and Pandey, 2024) developed an agricultural ontology for extracting relationships from text using Natural Language Processing, aiming to improve semantic understanding. Wang et al. (Wang, 2024) proposed a deep learning model based on the Transformer architecture, integrating a self-attention mechanism to capture long-distance dependencies in text for improved semantic understanding tasks.
@@
Despite these advancements, a significant gap remains in the current understanding of semantics: the lack of explicit attention to ethical definitions, such as popularity bias (Zhang et al., 2024b). While LLMs excel at generating semantically relevant recommendations, the training data often reflects existing popularity distributions, leading to the overexposure of popular items and the neglect of niche content (Nie et al., 2024). This imbalance can result in unfair outcomes for users with diverse preferences, negatively impacting both fairness and user satisfaction. Recent studies have started exploring fairness in RecLLM through fine-tuning and related techniques (Jiang et al., 2024; Hua et al., 2024), primarily focusing on group fairness (e.g., ensuring parity across demographic groups). However, the semantic understanding of popularity bias, particularly its impact on the interaction between items and users, remains largely unaddressed, leaving a significant gap in the current research landscape.

### v2

@@ 1. Introduction
As recommender systems strive to become more trustworthy, enhancing semantic understanding of core echical concepts, such as popularity bias, is essential for advancing cognitive capabilities of web applications (Zhang et al., 2025d; Gao et al., 2025). Popularity bias occurs when popular items are recommended too frequently, while other products that may deserve more attention are ignored (Abdollahpouri et al., 2019). This surface-level issue, however, masks a more complex interaction between user preferences, item rankings, and recommendation algorithms (Tang et al., 2024). In this context, semantic understanding refers to a model’s ability to grasp the deeper implications of popularity bias, rather than merely recognizing its surface-level features (Xu, 2025). In other words, a comprehensive semantic understanding of popularity bias requires recognizing not only the observable imbalance in recommendations but also the underlying causal mechanisms that give rise to it.
@@
Developing a deeper semantic understanding of popularity bias is key to addressing challenges such as the long-tail problem (Zhang et al., 2025c), product homogeneity (Lin et al., 2025), and personalized user satisfaction (Silva et al., 2021; Zhang et al., 2025a). Existing methods, such as PAAC (Cai et al., 2024), primarily focus on surface-level mitigation by increasing recommendation diversity or boosting exposure to non-popular items. Although terms like “diversity” and “debiasing” are frequently used in prompts, our investigation with Large Language Models (LLMs) reveals that such cues remain shallow representations of popularity bias (Klimashevskaia et al., 2024; Xiang et al., 2025). They fail to encode the causal semantics that link user behavior, item features, and system design to the emergence of bias.
@@ 2.1. Semantic Understanding for Recommender Systems
As web applications continue to evolve, achieving a deeper semantic understanding of complex definitions is crucial for enhancing the cognitive capabilities of models (Zhang et al., 2024b, 2025b). Bhattacharya and Pandey (Bhattacharya and Pandey, 2024) developed an agricultural ontology for extracting relationships from text using Natural Language Processing, aiming to improve semantic understanding. Wang et al. (Wang, 2024) proposed a deep learning model based on the Transformer architecture, integrating a self-attention mechanism to capture long-distance dependencies in text for improved semantic understanding tasks.
@@
Despite these advancements, a significant gap remains in the current understanding of semantics: the lack of explicit attention to ethical definitions, such as popularity bias (Zhang et al., 2024a). While LLMs excel at generating semantically relevant recommendations, the training data often reflects existing popularity distributions, leading to the overexposure of popular items and the neglect of niche content (Nie et al., 2024). This imbalance can result in unfair outcomes for users with diverse preferences, negatively impacting both fairness and user satisfaction. Recent studies have started exploring fairness in RecLLM through fine-tuning and related techniques (Jiang et al., 2024; Hua et al., 2024), primarily focusing on group fairness (e.g., ensuring parity across demographic groups). However, the semantic understanding of popularity bias, particularly its impact on the interaction between items and users, remains largely unaddressed, leaving a significant gap in the current research landscape.